OpenStudy (anonymous):

solve by the substitution method 5x-2y=-37 6x+57=y

OpenStudy (anonymous):

x=-11, y=-9

OpenStudy (anonymous):

5x-2y=-37 5x-2(6x+57)=-37 5x-12x-114=-37 -7x=77 so x= -11 6(-11)+57=y -66+57= y so y=-9 Hakkunamatata has right answer

OpenStudy (anonymous):

why is this better in openstudy, and not math?